2) The volume of the right cordate section of the brain was estimated from CT scans for patients with obsessive-compulsive diCan someone please explain this to me in details. thank you
Answer #1

Solution

Part (a)

Let

X = Volume of the right cordate section of the brain OCD patients,

Y = Volume of the right cordate section of the brain of healthy persons (control group)

Then, X ~ N(µ1, σ12) and Y ~ N(µ2, σ22)

Hypotheses:

Null H0: σ12 = σ22   Vs Alternative HA: σ22 > σ21 [since SD of control group > SD of OCD]

Test statistic:

F= s2/s12where s1and s2 are standard deviations based on n1 observations on X and n2 observations on Y.

Calculations

n1

10

n2

12

s1

0.08

s2

0.09

Fcal

1.265625

α

0.05*

Fcrit

3.102485

n1 - 1

9

n2 - 1

11

p-value

0.367619

* α = 0.05 is assumed

Distribution, Significance Level, α Critical Value and p-value

Under H0, F~ Fn1 – 1, n2 – 1

Critical value = upper α% point of Fn1 – 1, n2 – 1

p-value = P(Fn1 – 1, n2 – 1 > F(cal))

Using Excel Function: Statistical FINV and FDIST, these are found as shown in the above table.

Decision:

Since Fcal < Fcrit, or equivalently, since p-value > α. H0 is accepted.

Conclusion:

There is not sufficient evidence to suggest the two variances are unequal. Answer 1

Part (b)

Hypotheses:

Null: H0: µ1 = µ2 Vs Alternative: HA: µ1 µ2

Test Statistic:

t = (Xbar - Ybar)/[s√{(1/n1) + (1/n2)}] where

s2 = {(n1 – 1)s12 + (n2 – 1)s22}/(n1 + n2 – 2);

Xbar and Ybar are sample averages and s1,s2 are sample standard deviations based on n1 observations on X and n2 observations on Y respectively.

Calculations

Summary of Excel calculations is given below:

n1

10

n2

12

Xbar

0.34

Ybar

0.45

s1

0.08

s2

0.09

s^2

0.007335

s

0.085645

tcal

2.999659

α

0.05*

tcrit

2.085963

p-value

0.007081

* α = 0.05 is assumed

Distribution, Significance Level, α Critical Value and p-value:

Under H0, t ~ tn1 + n2 - 2. Hence, for level of significance α%, Critical Value = upper (α/2)% point of tn1 + n2 - 2 and p-value = P(tn1 + n2 - 2 > | tcal |).

Using Excel Function: Statistical TINV and TDIST, these are found to be as shown in the above table.

Decision:

Since | tcal | > tcrit, or equivalently since p-value < α, H0 is rejected.

Conclusion:

There is sufficient evidence to suggest that there is difference in the mean volume of this part of the brain between the two groups. Answer 2

Part (c)

100(1 - α) % Confidence Interval for (μ1 - μ2) is: (Xbar – Ybar) ± {(t2n – 2, α/2)(s)√{(1/n1) + (1/n2)}

substituting the values from Parts (a) and (b),

95% confidence interval for the difference in the mean volume of this part of the brain between the two groups is: [0.034, 0.186] Answer 3

Part (d)

Under Part (a), Type I error would occur when we conclude that the variances are different when in reality they are equal and Type II error would occur when we conclude that the variances are the same when in reality they are not.

Under Part (b), Type I error would occur when we conclude that the means are different between the two groups when in reality they are equal and Type II error would occur when we conclude that the means are equal between the two groups when in reality they are not. Answer

DONE